• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 22:02:58 on March 09, 2009

    Entry number 265511

    This entry is a followup to: 265477

    keyword=T6 trigger set to PSh.OR.Sh for cosmics run after HV change.

    The change in BB polarity caused a significant shift in the shower and pre-shower PMTs. Kalyan is setting the gains based on the cosmics data taken earlier.

    We'll soon take another cosmics run to record the 'after' spectra.